### v3

# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

Abstract
In this work we systematically review the recent advancements in code processing with language models, covering 50+ models, 30+ evaluation tasks, 170+ datasets, and 700 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also discuss code-specific features such as AST, CFG, and unit tests, along with their application in training code language models, and identify key challenges and potential future directions in this domain. We keep the survey open and updated on GitHub at https://github.com/codefuse-ai/Awesome-Code-LLM.
1 Introduction
Language modeling has advanced remarkably in recent years with the advent of pretrained Transformers (Vaswani et al., 2017) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2018). As large language models (LLMs) scaled to hundreds of billions of parameters and started to display early signs of artificial general intelligence (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023), their applications have also transcended text processing. Pioneered by Codex (Chen et al., 2021b), LLMs have achieved impressive results in code processing, giving rise to commercial products such as GitHub Copilot111https://github.com/features/copilot and open-source multi-billion code models such as StarCoder (Li et al., 2023g) and Code LLaMA (Rozière et al., 2023).
The application of pretrained Transformers in code processing, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for code in this work, covering 50+ models, 30+ downstream tasks, 170+ datasets, and 700 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
Related to our work, we are aware of several surveys on similar topics, with three works concurrent to us (Hou et al., 2023; Zheng et al., 2023b; She et al., 2023). These works, however, focus either on NLP side (Zan et al., 2023; Xu & Zhu, 2022) or SE side (Niu et al., 2023; Hou et al., 2023; Zheng et al., 2023b; She et al., 2023), and do not cover models, tasks, and challenges from the other side. For example, Zan et al. (2023) focus on LLMs for text-to-code generation, while giving little discussion of other evaluation tasks in software engineering community. Hou et al. (2023) and She et al. (2023), in contrast, comprehensively review works from SE venues such as ASE and ICSE, but cite only a handful of works from deep learning and NLP venues such as ACL, EMNLP, NeurIPS, and ICLR.
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
where level=0font=,
where level=1text width=5.5em,font=,
where level=2text width=4.5em,font=,
where level=3text width=1.8em,font=,
where level=4text width=5.5em,font=,
[Code LMs
[Raw LM
[LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2022), GPT-NeoX (Black et al., 2022),
BLOOM (Scao et al., 2022), LLaMA (Touvron et al., 2023a), GPT-4 (OpenAI, 2023),
LLaMA 2 (Touvron et al., 2023b), Phi-1.5 (Li et al., 2023h),
text width=27em,font=]
[Adapted LM
[Codex (Chen et al., 2021b), PaLM Coder (Chowdhery et al., 2022), Minerva (Lewkowycz et al., 2022),
PaLM 2* (Anil et al., 2023), Code LLaMA (Rozière et al., 2023),
text width=27em,font=]
[Specialized LM
[Encoder
[CuBERT (Kanade et al., 2020), CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021a),
SynCoBERT (Wang et al., 2021c), Code-MVP (Wang et al., 2022d), SCodeR (Li et al., 2022e),
text width=20.5em]
[Decoder
[CLM
[GPT-C (Svyatkovskiy et al., 2020), CodeGPT (Lu et al., 2021), PolyCoder (Xu et al., 2022),
CodeGen (Nijkamp et al., 2023b), PyCodeGPT (Zan et al., 2022), PanGu-Coder (Christopoulou et al., 2022),
CodeGeeX (Zheng et al., 2023a), Jam (Su et al., 2023),
Phi-1 (Gunasekar et al., 2023), CodeFuse (Di et al., 2023),
text width=16.8em]
[FIM
[InCoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023g),
text width=16.8em]
[UniLM
[CugLM (Liu et al., 2020), UniXcoder (Guo et al., 2022),
text width=20.5em]
[Encoder-Decoder
[PyMT5 (Clement et al., 2020),
T5-code (Mastropaolo et al., 2021),
DOBF (Lachaux et al., 2021),
PLBART (Ahmad et al., 2021),
CodeT5 (Wang et al., 2021e),
SPT-Code (Niu et al., 2022),
AlphaCode (Li et al., 2022g),
NatGen (Chakraborty et al., 2022a),
ERNIE-Code (Chai et al., 2023),
CodeT5+ (Wang et al., 2023d),
text width=20.5em]
[Code
Finetuning
[Instruction Finetuning
[WizardCoder (Luo et al., 2023), PanGu-Coder2 (Shen et al., 2023), OctoCoder (Muennighoff et al., 2023),
MFTCoder (Liu et al., 2023b),
text width=20.5em]
[Reinforcement Learning
[CompCoder (Wang et al., 2022c), CodeRL (Le et al., 2022), PPOCoder (Shojaee et al., 2023), RLTF (Liu et al., 2023d),
text width=20.5em]
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2023a; Luo et al., 2023), infilling objectives (Tay et al., 2023; Li et al., 2023g; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023h), architectural improvements (Shazeer, 2019; Su et al., 2021; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities.
The rest of this work is organized following the taxonomy presented in Figure 1. In Section 2 we first provide the preliminaries of language modeling and Transformer models, and then in Section 3 we contextualize the evaluation of language models for code, highlighting the historical transition from various code understanding tasks to more practical text-to-code generation tasks. In Section 4 we discuss the plethora of LLMs that have demonstrated coding ability, and then in Section 5 we review the specialized and often smaller models by their architecture, with special attention on the recent application of infilling objectives, instruction tuning, reinforcement learning, and engineering improvements. Then, in Section 6, we discuss unique features of code that are not available to natural languages but have been utilized to aid code processing. In Section 7, we review the most recent integration between LLMs and software development, before finally concluding this work in Section 8 and highlighting the current challenges in code processing.
2 Background
In this section, we briefly review the preliminaries of Transformer-based language modeling, including common objectives for unidirectional and bidirectional models, as well as some popular models and designs in NLP.
2.1 Causal Language Modeling
Unidirectional language models (also known as causal language models222The training objective of such language models is Causal Language Modeling (CLM), but also referred to as Next Token Prediction.) factor the probability of a sentence into the product of each token’s conditional probability with the chain rule. A piece of input text consisting of tokens is modeled as
where is a shorthand for tokens before in the input, and is the parameters of the model. With Transformer decoders such as GPT (Radford et al., 2018; 2019; Brown et al., 2020) and LLaMA (Touvron et al., 2023a; b), the conditional probability in (1) is modeled by adding an attention mask to the attention matrix of each Transformer block, ensuring that can only attend to previous tokens. During training, the cross entropy loss on all tokens in the input is calculated in parallel, while at inference time each new token is generated autoregressively. For further details about the Transformer architecture we refer to Vaswani et al. (2017).
2.2 Masked Language Modeling
Unlike causal language models, bidirectional language models are trained to acquire a better contextual representation of text rather than generating text autoregressively. In the vanilla Transformer, the encoder part is allowed to attend to a token’s left as well as right context for this purpose. BERT (Devlin et al., 2019) takes one step further and pretrains only a Transformer encoder. A set of randomly chosen tokens in the input are replaced by a special token [MASK] to obtain a noisy input , for example 333Both [CLS] and [EOS] are artificial tokens added to the input text. [CLS] is added at the beginning and its representation is used for sentence classification, while [EOS] indicates end of sentence. The original BERT also uses another special token [SEP], which is not in common use in LLMs, and we refer to Devlin et al. (2019) for details., and the model is trained to recover the original tokens by maximizing
While this objective requires the model to have a deep understanding of the input text to reconstruct it, it suffers from low training efficiency, since only a small set of tokens (usually 15%) are masked (and thus “trained on”). To address this issue, Clark et al. (2020) propose ELECTRA, which is trained to discriminate whether or not each token in the input has been replaced by a BERT-like model instead.
2.3 Denoising Objectives
GPT-style causal LM and BERT-style bidirectional LM each has its own strengths and weaknesses. While GPT can be used for autoregressive generation, it lacks a bidirectional representation of input text, and is thus unsuitable for sequence-to-sequence (seq2seq) generation tasks such as translation and summarization. BERT, on the other hand, can produce bidirectional representations, but is pretrained only for mask filling, not generation.
The vanilla Transformer encoder-decoder architecture combines the respective advantages of GPT and BERT. T5 (Raffel et al., 2020) is such a model pretrained with span corruption, which can be regarded as a variant of MLM. During pretraining, spans of text in the input are replaced with sentinel tokens, which plays the same role as [MASK] in BERT. The noisy input is first processed by the encoder with bidirectional attention, and the masked spans are then generated autoregressively by the decoder. Formally, if spans are sampled for corruption in input , the noisy input is then constructed by replacing each span with a special token <extra_id_i>, for , and the target is constructed by concatenating all spans prepended with corresponding sentinels: [<extra_id_1>, span, , <extra_id_k>, span]. The model is then trained with a standard seq2seq objective, by maximizing
Lester et al. (2021) show that models pretrained with such objectives can be adapted for autoregressive language modeling with extra pretraining using the prefix language modeling objective, i.e. spliting the text into two parts, processing the first part with encoder and generating the second part with decoder.
Tay et al. (2023) argue that span corruption is also closely related to CLM, since one can mask out the whole input text as a single span and train the decoder to generate it autoregressively. Inspired by this relation, they propose UL2, which is the combination of many span corruption objectives that differ in corruption rate and span length. Applying it to both encoder-decoder models and decoder-only models, they find that encoder-decoder models perform better under the same computation budget constraint. Other researches have also found that such encoder-decoder models generally perform better than causal decoder-only models (Wang et al., 2022b; Soltan et al., 2022).
2.4 Auxiliary Objectives
Language modeling objectives, such as previously discussed CLM and MLM, mainly train the model to capture token-level information and are ineffective at modeling document structures. Thus, auxiliary objectives are often added to help the models learn such global information. BERT is pretrained with next sentence prediction (NSP) along with MLM, which is formulated as a binary classification task to predict whether two segments in the input are neighboring in the original corpus. Lan et al. (2020) propose a more challenging sentence-order prediction (SOP) task, where the negative samples are constructed by swapping the order of two neighboring sentences instead of sampling a random sentence from other documents.
Relatedly, Raffel et al. (2020) mix supervised downstream samples such as GLUE (Wang et al., 2018a) into T5’s pretraining dataset to conduct multi-task pretraining. However, it is worth noting that since they unify all tasks into a text-to-text format, the training objective is the same for their self-supervised pretraining and supervised downstream tasks, i.e. conditional generation of the target sequence.
2.5 Implementation Design
While most researches on pretraining language models have focused on designing novel training objectives, low-level implementation of the Transformer architecture itself is also being continuously improved over the years in pursuit of stability, performance, and efficiency, as shown in Figure 2.
The original Transformer block proposed by Vaswani et al. (2017) is formulated as
where is the layer’s input; is the layer’s output; “Attention” is the self-attention sublayer; “FFN” is the feed-forward sublayer, and “LN” is layer normalization (Ba et al., 2016).
GPT-2 (Radford et al., 2019) moves layer normalization to the input of each Transformer sub-block to stabilize training:
and such pre-norm has since become a standard practice in Transformer decoders.
GPT-J (Wang & Komatsuzaki, 2021) modifies the Transformer block to compute FFN sub-layer and self-attention sub-layer in parallel to increase computation throughput:
and Chowdhery et al. (2022) observes limited performance degradation when applying this design to larger models.
PaLM (Chowdhery et al., 2022) introduces Rotary Position Embedding (RoPE) and Multi-Query Attention (MQA) into LLMs. RoPE (Su et al., 2021) multiplies the keys and queries of each self-attention layer by a position-dependent rotation matrix to inject position information, and is later shown to enable position interpolation for processing of longer sequences (Chen et al., 2023b; Rozière et al., 2023). Alternative to RoPE, Press et al. (2022) propose ALiBi, which directly attenuates the attention scores according to the relative position between key and query. This position embedding scheme is later adopted by BLOOM (Scao et al., 2022).
Apart from position embeddings, another issue in Transformer that has long troubled researchers is the fact that the complexity of self-attention scales quadratically with the input sequence length. Some works such as Reformer (Kitaev et al., 2020), Linformer (Wang et al., 2020c), Performer (Choromanski et al., 2021) and cosFormer (Qin et al., 2022b) use approximate attention to reduce this complexity, but they mostly come at the cost of degraded performance. Other works tackle this issue from an engineering point-of-view. MQA (Shazeer, 2019) shares the same set of keys and values across all attention heads to optimize memory-to-arithmetic ratio and significantly improves inference speed at small costs of model performance. Its variant Grouped-Query Attention (GQA, Ainslie et al., 2023) takes a middle-ground approach by dividing attention heads into groups and sharing the same set of keys/values within each group. Orthogonally, Dao et al. (2022) introduce FlashAttention, which is an exact but improved implementation of self-attention that optimizes IO operations on the accelerating device via tiling to improve memory efficiency.
Another important technique for improving LLMs’ training efficiency is mixed precision training (Micikevicius et al., 2018), which stores model weights and activations in lower precision to save memory consumption. Early works use FP16 format for this low precision, while BF16 has now become more popular (Chowdhery et al., 2022; Scao et al., 2022), as it can represent the same range of floating point numbers as FP32, thus preventing numerical overflowing and underflowing during training.
3 Evaluation of Language Models for Code
Over the past decade, various evaluation tasks have been proposed by the software engineering community to evaluate code models. CodeXGLUE (Lu et al., 2021) consolidates most of such tasks into a single benchmark covering code understanding tasks such as clone detection, defect detection, and sequence-to-sequence generation tasks such as code repair, code translation, program synthesis, and code summarization. After Chen et al. (2021b) introduced HumanEval and Codex, however, text-to-code synthesis was brought into the spotlight in the NLP community and has since become a standard task for evaluating LLMs (Figure 3). Accordingly, we first briefly introduce each of the traditional tasks and the application of pretrained language models in them in Section 3.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in Section 3.2 and investigate program synthesis in more detail in Section 3.3. Lastly, we also discuss the latest trend of repository-level evaluation in Section 3.4. In Appendix A, we list benchmarks for each downstream task.
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=6em,font=,
where level=4text width=8em,font=,
where level=5text width=13em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation
[Text-to-Code
[Code
Retrieval
Lu et al. (2015), CodeHow (Lv et al., 2015),
RACS (Li et al., 2016), CodeMatcher (Liu et al., 2022a)
CODE-NN (Iyer et al., 2016), DeepCS (Gu et al., 2018),
Yin et al. (2018), NCS (Sachdev et al., 2018),
UNIF (Cambronero et al., 2019), HECS (Li et al., 2020b),
CARLCS-CNN (Shuai et al., 2020), DGMS (Ling et al., 2021),
NJACS (Hu et al., 2020),
TabCS (Xu et al., 2021), GraphSearchNet (Liu et al., 2023h),
TranCS (Sun et al., 2022)
TranS (Wang et al., 2020e), Heyman & Cutsem (2020),
Corder (Bui et al., 2021b), SAN-CS (Fang et al., 2021),
SST (Gu et al., 2021),
MuCoS (Du et al., 2021), MEM (Salza et al., 2023),
CDCS (Chai et al., 2022), CodeRetriever (Li et al., 2022d),
CoCoSoDa (Shi et al., 2023b), Li et al. (2022a),
CCT-LM (Sorokin et al., 2023)
[Code
Synthesis
Euphony (Lee et al., 2018), Neo (Feng et al., 2018)
LPN (Ling et al., 2016), NSPS (Parisotto et al., 2017),
DeepCoder (Balog et al., 2017), RobustFill (Devlin et al., 2017a),
(Yin & Neubig, 2017), ASN (Rabinovich et al., 2017),
NGDS (Kalyan et al., 2018), Bunel et al. (2018),
ReCode (Hayati et al., 2018), AutoPandas (Bavishi et al., 2019),
Wei et al. (2019), PlotCoder (Chen et al., 2021c),
Huang et al. (2022a)
TreeGen (Sun et al., 2020), REDCODER (Parvez et al., 2021),
Jigsaw (Jain et al., 2022), JuPyT5 (Chandel et al., 2022),
CodeT (Chen et al., 2023a), TiCoder (Lahiri et al., 2022),
AceCoder (Li et al., 2023d), Self-Debugging (Chen et al., 2023c),
ClarifyGPT (Mu et al., 2023)
[Text-to-SQL
Seq2SQL (Zhong et al., 2017), SQLNet (Xu et al., 2017),
Suhr et al. (2018),
TypeSQL (Yu et al., 2018a), Coarse2Fine (Dong & Lapata, 2018),
Finegan-Dollak et al. (2018), SyntaxSQLNet (Yu et al., 2018b),
GNN (Bogin et al., 2019), TREQS (Wang et al., 2020b)
SQLova (Hwang et al., 2019), IRNet (Guo et al., 2019),
Zhang et al. (2019b), RAT-SQL (Wang et al., 2020a),
Bertrand-DR (Kelkar et al., 2020), RYANSQL (Choi et al., 2021),
TaBERT (Yin et al., 2020),
Photon (Zeng et al., 2020), HydraNet (Lyu et al., 2020),
GAZP (Zhong et al., 2020), GraPPa (Yu et al., 2021),
SmBoP (Rubin & Berant, 2021), NQG-T5 (Shaw et al., 2021),
StruG (Deng et al., 2021), SLSQL (Lei et al., 2020),
GAP (Shi et al., 2021), Wang et al. (2021a),
GP (Zhao et al., 2021),
LGESQL (Cao et al., 2021), Picard (Scholak et al., 2021),
H-NeurSyn (Yang et al., 2021),
UnifiedSKG (Xie et al., 2022b), CodexDB (Trummer, 2022),
T5QL (Arcadinho et al., 2022), TKK (Gao et al., 2022),
Graphix-T5 (Li et al., 2023e),
RESDSQL (Li et al., 2023b), Liu et al. (2023a),
Self-Debugging (Chen et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023),
Chang & Fosler-Lussier (2023),
Nan et al. (2023), SQL-PaLM (Sun et al., 2023a),
Guo et al. (2023a), DAIL-SQL (Gao et al., 2023a)
[Math Programming
MathQA-Python (Austin et al., 2021), Drori & Verma (2021),
Drori et al. (2022),
PAL (Gao et al., 2023b), PoT (Chen et al., 2022),
He-Yueya et al. (2023),
GSM8K-Python (Chowdhery et al., 2022; Wang et al., 2023d)
[Code-to-Code]
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
3.1 Downstream Tasks of Code Processing
Following the custom in software engineering, we categorize the evaluation tasks for code according to their input/output modality, and break down these tasks into five families: text-to-code, code-to-code, code-to-text, code-to-pattern, and text-to-text. We note that this taxonomy is interleaved with the understanding-generation dichotomy in NLP, since each category may contain both understanding and generation tasks, as discussed in Section 3.1.6.
3.1.1 Text-to-Code
Text-to-code tasks take text as input, and output code. Related methods are listed in Figure 4
- Code retrieval aims to retrieve relevant code given natural language queries, or to mine parallel text-code pairs from an unannotated corpus. This task is usually performed by computing a similarity metric between the embedding of query and candidate code, and the contextual embeddings produced by bidirectional language models - such as BERT - has proven to be extremely helpful. Grazia & Pradel (2023) and Xie et al. (2023a) provide comprehensive reviews on this topic.
- Code synthesis aims to generate code (usually a function or a method) given a natural language description. This task can be viewed as an updated version of code retrieval using generative models instead of retrieval models. Statistical machine translation (SMT) and neural machine translation (NMT) models have been widely adopted for this task, often with enhanced decoders that leverage the unique grammatical rules of programming languages (Yin & Neubig, 2017; Rabinovich et al., 2017). Pretrained language models based on Transformer architecture, however, changed the game by directly generating the source code in the autoregressive language modeling style, even without task-specific finetuning (Chen et al., 2021b). We discuss this task in more detail in Section 3.3.
- Text-to-SQL is a special (and arguably easier) case of code synthesis, where the model is tasked to generate SQL commands from natural language queries. It has been a topic of special interest due to SQL’s structured nature (when compared with general-purpose languages such as Python and C) and wide application in data management. We refer to Kumar et al. (2022); Deng et al. (2022a); Qin et al. (2022a); Katsogiannis-Meimarakis & Koutrika (2023) for surveys on this topic.
- Math programming is also a special case of code synthesis, where a language model is required to solve mathematical reasoning problems via generating code that will be executed by external interpreters. This task abstracts the reasoning process from numerical calculations, and is thus of special interest in evaluating LLMs.
3.1.2 Code-to-Code
Code-to-code tasks take code as input, and output code. Related methods are listed in Figure 5 and 6
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=8em,font=,
where level=4text width=8em,font=,
where level=5text width=11em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation
[Text-to-Code
[Code-to-Code
[Code Search
FaCoY (Kim et al., 2018), Aroma (Luan et al., 2019),
COSAL (Mathew & Stolee, 2021)
Corder (Bui et al., 2021b), Cosco (Saieva et al., 2023)
[Code
Completion
BMN (Bruch et al., 2009), MSE (Hindle et al., 2012), Naturalize (Allamanis et al., 2014),
Cache LM (Tu et al., 2014), DeepSyn (Raychev et al., 2016b), PHOG (Bielik et al., 2016),
Deep3 (Raychev et al., 2016a), Hellendoorn & Devanbu (2017)
Raychev et al. (2014), White et al. (2015),
Pointer Mixture Network (Li et al., 2018a), Parvez et al. (2018),
SLM (Alon et al., 2020), Pythia (Svyatkovskiy et al., 2019), NLM (Karampatsis et al., 2020)
IntelliCode (Svyatkovskiy et al., 2020), CugLM (Liu et al., 2020),
eWASH (Clement et al., 2021),
LongCoder (Guo et al., 2023b)
[Code
Translation
lpSMT (Nguyen et al., 2013), Karaivanov et al. (2014),
mppSMT (Nguyen et al., 2015)
Tree2Tree (Chen et al., 2018), Grammar Tree2Tree (Drissi et al., 2018)
TransCoder (Rozière et al., 2020),
TransCoder-ST (Rozière et al., 2022),
TransCoder-IR (Szafraniec et al., 2023),
BabelTower (Wen et al., 2022),
SDA-Trans (Liu et al., 2023c),
Self-Debugging (Chen et al., 2023c),
CoTran (Jana et al., 2023),
MuST (Zhu et al., 2022b),
Pan et al. (2023), Tang et al. (2023)
[API
Mining
RASH (Zhang et al., 2018), RAPIM (Alrubaye et al., 2020)
Nguyen et al. (2016), DeepAM (Gu et al., 2017),
JV2CS (Phan et al., 2017), Lu et al. (2017),
TL-CodeSum (Hu et al., 2018b),
BIKER (Huang et al., 2018), Chen et al. (2021a),
SAR (Bui et al., 2019)
HaPiM (Zhou et al., 2023b)
[Cloze Test
CodeBERT (Feng et al., 2020),
Lu et al. (2021),
Puri et al. (2021),
JavaBERT (de Sousa & Hasselbring, 2021)
[Code
Repair
Prophet (Long & Rinard, 2016), TBar (Liu et al., 2019b),
Refactory (Hu et al., 2019), PyTER (Oh & Oh, 2022)
sk_p (Pu et al., 2016),
DeepFix (Gupta et al., 2017), SSC (Devlin et al., 2017b),
SynFix (Bhatia et al., 2018),
Codit (Chakraborty et al., 2022b), Tufano et al. (2019b),
Sequencer (Chen et al., 2021e), Tufano et al. (2019a),
Vasic et al. (2019), DrRepair (Yasunaga & Liang, 2020),
CoCoNuT (Lutellier et al., 2020), DLFix (Li et al., 2020c),
Review4Repair (Huq et al., 2022),
DEAR (Li et al., 2022f)
CURE (Jiang et al., 2021a),
DeepDebug (Drain et al., 2021), BIFI (Yasunaga & Liang, 2021),
Recoder (Zhu et al., 2021), TFix (Berabi et al., 2021),
Modit (Chakraborty & Ray, 2021), Fan et al. (2023),
AlphaRepair (Xia & Zhang, 2022), RING (Joshi et al., 2023),
Xia et al. (2023), VulRepair (Fu et al., 2022),
CodeT5-DLR (Bui et al., 2022),
Conversational APR (Xia & Zhang, 2023), Jiang et al. (2023a),
Paul et al. (2023), Cao et al. (2023),
TypeFix (Peng et al., 2023b)
[Code
Infilling
InCoder (Fried et al., 2023), FIM (Bavarian et al., 2022), SantaCoder (Allal et al., 2023),
StarCoder (Li et al., 2023g), Code LLaMA (Rozière et al., 2023)
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=10em,font=,
where level=4text width=6em,font=,
where level=5text width=11em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation
[Text-to-Code
[Code-to-Code
[Deobfuscation
DeGuard (Bichsel et al., 2016), Autonym (Vasilescu et al., 2017),
Debin (He et al., 2018), JSNeat (Tran et al., 2019)
DIRE (Lacomis et al., 2019), Artuso et al. (2021)
VarBERT (Banerjee et al., 2021), DIRECT (Nitin et al., 2021),
SLaDe (Armengol-Estapé et al., 2023), LmPa (Xu et al., 2023b)
[Unit Test Generation
EvoSuite (Fraser & Arcuri, 2011), EvoSuiteR (Shamshiri, 2015),
DynaMOSA (Panichella et al., 2018), LambdaTester (Selakovic et al., 2018),
TSE (Shimmi & Rahimi, 2022), Nessie (Arteca et al., 2022)
AthenaTest (Tufano et al., 2021a), FSLM (Bareiß et al., 2022),
TestPilot (Schäfer et al., 2023), A3Test (Alagarsamy et al., 2023),
TeCo (Nie et al., 2023), CodaMosa (Lemieux et al., 2023),
ChatTester (Yuan et al., 2023b), ChatUniTest (Xie et al., 2023b)
[Assertion Generation
MeMo (Blasi et al., 2021)
Atlas (Watson et al., 2020)
Tufano et al. (2022a), TOGA (Dinella et al., 2022),
FSLM (Bareiß et al., 2022), Yuan et al. (2023a)
[Mutant
Generation
Major (Just, 2014), LAVA (Dolan-Gavitt et al., 2016),
PIT (Coles et al., 2016), Allamanis et al. (2016a),
EvilCoder (Pewny & Holz, 2016), wild-caught mutants (Brown et al., 2017),
Apocalypse (Roy et al., 2018), Bug-Injector (Kashyap et al., 2019),
IBIR (Khanfir et al., 2023b), PBMT (Bartocci et al., 2023)
Tufano et al. (2019c), DeepMutation (Tufano et al., 2020),
SemSeed (Patra & Pradel, 2021)
BERT (Degiovanni & Papadakis, 2022; Khanfir et al., 2023a),
FSLM (Bareiß et al., 2022)
[Fuzzing
SymFuzz (Cha et al., 2015), AFLFast (Böhme et al., 2019),
FairFuzz (Lemieux & Sen, 2018), AFLGo (Böhme et al., 2017),
Angora (Chen & Chen, 2018), TensorFuzz (Odena et al., 2019),
Audee (Guo et al., 2020), LEMON (Wang et al., 2020f),
DocTer (Xie et al., 2022a), FreeFuzz (Wei et al., 2022a),
SpecFuzzer (Molina et al., 2022), Muffin (Gu et al., 2022),
DeepREL (Deng et al., 2022b), NNSmith (Liu et al., 2023e),
Fuzz (Yang et al., 2023c)
NEUZZ (She et al., 2019), MTFuzz (She et al., 2020),
PreFuzz (Wu et al., 2022)
TitanFuzz (Deng et al., 2023),
WhiteFox (Yang et al., 2023b)
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
- Code search is a task similar to code retrieval, and differs from the later only in that the input is an existing code snippet, often in a different programming language from the target. Code search can be reformulated as finding clones of the query in the pool of targets, and is thus equivalent to clone detection to some extent.
- Code completion aims to complete a piece of code given its prefix, and remains to date one of the most popular applications of code language models in IDEs. This is essentially language modeling applied to code, and related technologies have been progressively introduced: n-gram, RNN, and Transformer. However, due to the structured nature of programming languages, many early works found grammar-aided statistical models to perform better (Bielik et al., 2016; Hellendoorn & Devanbu, 2017), and neural models only became dominant after 2018 (see Figure 5 for an intuitive overview.)
- Code translation aims to translate a piece of code (usually a function or method) into another programming language. The relation between code translation and cross-lingual code search is similar to the one between code synthesis and text-to-code retrieval, and SMT/MNT models have also been widely applied to this task. Code translation is an important technique in migrating old projects written in obsolete languages. However, we are yet to witness such applications, as the context window of even the most powerful language models are quite limited in the face of such projects. Malyala et al. (2023) provide a short survey on this task from the SE perspective.
- API mining refers to the process of finding similar APIs in different libraries, potentially in different programming languages. This task is closely related to API migration - a special case of code translation. API mining is traditionally tackled by computing similarity metrics between source and target APIs using information retrieval models, but as generative models become ever more capable, it is also worth exploring to directly generate the target API as a sequence-to-sequence task. Another closely related task is idiom mining (Allamanis & Sutton, 2014), where the objective is to discover commonly used code patterns, which exposes the potential need of new APIs (Sivaraman et al., 2022).
- Cloze test is a recently proposed task for code understanding, after the rise of BERT-style pretraining. Due to the unique semantics of programming languages, several keywords are often selected for this test, such as min and max (Feng et al., 2020).
- Code repair, also known as bug fix, aims to fix a piece of buggy code. Like code translation, it is a traditional sequence-to-sequence generation task, and surveys are abundant on this topic (Gazzola et al., 2018; Monperrus, 2018; Zhong et al., 2022; Zhang et al., 2023c; Huang et al., 2023a). A generalization of code repair is code edit or code change, where the task is to model and predict code changes (such as Git commits), which are not necessarily about bug fixing.
- Code infilling is another recently proposed task, after fill-in-the-middle pretraining (Bavarian et al., 2022) became popular. It is a generalization of code completion, where not only the left context, but also the right context is given. However, it differs from cloze test in that the target of cloze test is only one token, while the target of code infilling can be an entire line or even multiple lines, which requires a decoder to generate autoregressively.
- Obfuscation refers to the process of renaming identifiers (e.g. variables, methods, and classes), for example to generic names like var_1, var_2 or x, y. It is an important technique in virus detection, intellectual property protection, and code size reduction (Collberg & Thomborson, 2002; Murad et al., 2010; Vasilescu et al., 2017). Deobfuscation refers to the reverse process, where meaningful identifier names are recovered from obfuscated programs. Obfuscation can be easily achieved statically, but deobfuscation has been a subject of more interest in recent years. It plays a significant role in decompiling or reverse engineering, and has also been adopted as a pretraining objective for code language models (Lachaux et al., 2021; Ding et al., 2022a; Liu et al., 2022d).
- Unit test generation aims to generate unit tests for a given program. Prior to the rise of Codex and other code LLMs, almost all works in this area employed non-neural methods (see Figure 6). In the age of LLMs, however, this task is ever more important, as researches have shown that the current unit tests for evaluating LLMs’ program synthesis capability may be insufficient (Liu et al., 2023f).
- Assertion generation is a subtask of unit testing. Given a program and a partial unit test, this task aims to generate assertions (also known as oracles in software engineering) within the unit test. This task has generally went unnoticed by the NLP community, as the program synthesis task used for evaluating LLMs often concern standalone, competition-style methods, for which the simple assertion of the equality between program output and expected answer suffices.
- Mutant generation aims to generate mutants of a given program for the purpose of mutation testing, and relates closely to unit test generation and assertion generation. A mutant that is not detected by a given set of unit tests and assertions indicates that either additional test cases or better assertions are required (Fraser & Arcuri, 2011). Recently, masking out tokens in the source code and sampling them from the output of a masked language model has become a common method for this task. Ojdanic et al. (2021; 2023) give empirical comparisons between different mutation methods.
- Fuzzing is another software testing task, where the objective is to generate a large set of inputs covering as many corner cases as possible. While many recent works on fuzzing target deep learning libraries, few have utilized language models to conduct this process (see Figure 6).
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=7em,font=,
where level=4text width=9em,font=,
where level=5text width=11em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation
[Text-to-Code]
[Code-to-Code]
[Code-to-Text
[Code Summarization
CloCom (Wong et al., 2015)
CODE-NN (Iyer et al., 2016), DeepCom (Hu et al., 2018a),
TL-CodeSum (Hu et al., 2018b), Code2Seq (Alon et al., 2019a),
Fernandes et al. (2019), Wan et al. (2018),
AST-AttendGRU (LeClair et al., 2019), Wei et al. (2019),
Haque et al. (2020), DMACOS (Xie et al., 2021),
Bansal et al. (2021),
CoCoSUM (Wang et al., 2021d),
MLCS (Zhou et al., 2023c)
TranS (Wang et al., 2020e), Ahmad et al. (2020),
Corder (Bui et al., 2021b), SiT (Wu et al., 2021),
SG-Trans (Gao et al., 2023c),
Codex-D (Chen et al., 2021b),
M2TS (Gao & Lyu, 2022), AST-Trans (Tang et al., 2022),
GTrans (Kuang et al., 2022), CoSS (Shi et al., 2023a),
Sun et al. (2023b), Yuan et al. (2023a),
use-seq (Su & McMillan, 2023a),
Su & McMillan (2023b)
[Code
Review
Review Bot (Balachandran, 2013), Tricorder (Sadowski et al., 2015),
ClusterChanges (Barnett et al., 2015), CommentFinder (Hong et al., 2022)
DCR (Gupta & Sundaresan, 2018), DeepReview (Li et al., 2019a),
DACE (Shi et al., 2019), CORE (Siow et al., 2020)
Tufano et al. (2021b), Hellendoorn et al. (2021),
Tufano et al. (2022b), CodeReviewer (Li et al., 2022i),
AUGER (Li et al., 2022b), crBERT (Yin et al., 2023),
LLaMA-Reviewer (Lu et al., 2023)
[Identifier Prediction
JSNice (Raychev et al., 2015), Pigeon (Alon et al., 2018),
HeMa (Jiang et al., 2019)
Allamanis et al. (2015), Allamanis et al. (2016b),
GGNN (Allamanis et al., 2018), Code2Vec (Alon et al., 2019b),
Code2Seq (Alon et al., 2019a), Fernandes et al. (2019),
HIER (Xu et al., 2019b), Liu et al. (2019a),
MNire (Nguyen et al., 2020), DeepName (Li et al., 2021b),
DMACOS (Xie et al., 2021), NamPat (Yang et al., 2022)
GTNM (Liu et al., 2022b), GTrans (Kuang et al., 2022),
Mario (Wang et al., 2023b)
[Commit Message Generation
ChangeScribe (Cortes-Coy et al., 2014; Vásquez et al., 2015), Jiang & McMillan (2017),
NNGen (Liu et al., 2018), ChangeDoc (Huang et al., 2020)
Loyola et al. (2017), Jiang et al. (2017),
Loyola et al. (2018), PtrGNCMsg (Liu et al., 2019c),
CoDiSum (Xu et al., 2019a), ATOM (Liu et al., 2022c),
CC2Vec (Hoang et al., 2020), CoRec (Wang et al., 2021b)
CoreGen (Nie et al., 2021), CommitBERT (Jung, 2021),
RACE (Shi et al., 2022), FIRA (Dong et al., 2022),
COME (He et al., 2023a), ExGroFi (Wang et al., 2023a)
[Code-to-Pattern
[Text-to-Text
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=6em,font=,
where level=4text width=10em,font=,
where level=5text width=11em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation
[Text-to-Code]
[Code-to-Code]
[Code-to-Text]
[Code-to-Pattern
[Type
Prediction
JSNice (Raychev et al., 2015), TypeDevil (Pradel et al., 2015),
Xu et al. (2016), Pigeon (Alon et al., 2018),
Typpete (Hassan et al., 2018)
DeepTyper (Hellendoorn et al., 2018), Jangda & Anand (2019),
NL2Type (Malik et al., 2019), DLTPy (Boone et al., 2019),
TypeWriter (Pradel et al., 2020), OptTyper (Pandi et al., 2020),
Typilus (Allamanis et al., 2020), LambdaNet (Wei et al., 2020),
Type4Py (Mir et al., 2022), HiTyper (Peng et al., 2022),
PYInfer (Cui et al., 2021)
TypeBert (Jesse et al., 2021), TypeWeaver (Yee & Guha, 2023),
TypeT5 (Wei et al., 2023), OpenTau (Cassano et al., 2023b),
TypeGen (Peng et al., 2023c)
[Defect
Detection
Ray et al. (2016), Bugram (Wang et al., 2016a),
NAR-Miner (Bian et al., 2018), (Pearce et al., 2022)
Wang et al. (2016b), VulDeePecker (Li et al., 2018c),
Lin et al. (2018a),
DeepBugs (Pradel & Sen, 2018), Russell et al. (2018),
SySeVR (Li et al., 2018b),
Devign (Zhou et al., 2019), Li et al. (2019c),
Lin et al. (2021),
VulDeeLocator (Li et al., 2022h), VulDeePecker (Zou et al., 2021),
ReVeal (Chakraborty et al., 2022c), BugLab (Allamanis et al., 2021),
IVDetect (Li et al., 2021a),
ReGVD (Nguyen et al., 2022)
GREAT (Hellendoorn et al., 2020), VulBERTa (Hanif & Maffeis, 2022),
LineVul (Fu & Tantithamthavorn, 2022),
DeepDevVuln (Chan et al., 2023), Yuan et al. (2023a),
(Zhang et al., 2023a), CausalVul (Rahman et al., 2023),
Gao et al. (2023d)
[Clone
Detection
Deckard (Jiang et al., 2007), SourcererCC (Sajnani et al., 2016),
CCAligner (Wang et al., 2018b), LVMapper (Wu et al., 2020),
SAGA (Li et al., 2020a),
NIL (Nakagawa et al., 2021)
White et al. (2016), CDLH (Wei & Li, 2017),
Oreo (Saini et al., 2018), DeepSim (Zhao & Huang, 2018),
ASTNN (Zhang et al., 2019a),
TBCCD (Yu et al., 2019a), Perez & Chiba (2019)
CLCDSA (Nafi et al., 2019),
FA-AST (Wang et al., 2020d),
Fang et al. (2020), Yahya & Kim (2022)
Khajezade et al. (2022),
SSCD (Chochlov et al., 2022),
CCT-LM (Sorokin et al., 2023),
Dou et al. (2023),
Yuan et al. (2023a),
ZC (Li et al., 2023c)
[Code
Reasoning
[MMLU (Hendrycks et al., 2021b),
CodeQA (Liu & Wan, 2021),
CS1QA (Lee et al., 2022),
C-Eval (Huang et al., 2023b),
CMMLU (Li et al., 2023a),
CodeApex (Fu et al., 2023)
[Code/Author Classification
SCAP (Frantzeskou et al., 2011), Mahbub et al. (2022)
[Bandara & Wijayarathna (2013),
TBCNN (Mou et al., 2016), inst2vec (Ben-Nun et al., 2018),
DL-CAIS (Abuhamad et al., 2018),
ASTNN (Zhang et al., 2019a), InferCode (Bui et al., 2021a)
Puri et al. (2021), Wang et al. (2022a)
[Machine Code
Detection
DetectGPT4Code (Yang et al., 2023e)
[Text-to-Text
[Document Translation
[Lu et al. (2021)]
[Log
Analysis
[Spell (Du & Li, 2016), Drain (He et al., 2017),
Logram (Dai et al., 2022), SPINE (Wang et al., 2022e)
DeepLog (Du et al., 2017),
Lu et al. (2018),
LogAnomaly (Meng et al., 2019),
LogRobust (Zhang et al., 2019c),
LogDTL (Nguyen et al., 2021),
UniParser (Liu et al., 2022e)
NuLog (Nedelkoski et al., 2020a),
Logsy (Nedelkoski et al., 2020b),
LogBERT (Guo et al., 2021b),
NeuralLog (Le & Zhang, 2021), LogStamp (Tao et al., 2022), LogPPT (Le & Zhang, 2023a),
Le & Zhang (2023b), Semlog (Yu et al., 2023b),
LogPrompt (Liu et al., 2023k), LogGPT (Qi et al., 2023),
Mudgal & Wouhaybi (2023), LLMParser (Jiang et al., 2023b)
3.1.3 Code-to-Text
Code-to-text tasks take code as input, and output text. Related methods are listed in Figure 7.
- Code summarization, also referred to as docstring generation, aims to generate a natural language description for a given piece of code (often a function or method). This is the opposite of code synthesis, and SMT/NMT techniques have been likewise applied. Zhang et al. (2022) provide a survey on this topic.
- Code review aims to automate the process of peer code review, and includes many subtasks, such as review necessity prediction, review comment generation, code refinement, and review decision prediction.
- Identifier prediction is the task of predicting identifier names in the code. As these names are deemed to contain important semantic information, this task has been utilized for code summarization (Allamanis et al., 2016b), as well as pretraining code models (Wang et al., 2021e; Niu et al., 2022). A special case of identifier prediction is method name prediction.
- Commit message generation aims to automatically generate commit messages for code changes. This task takes the code before and after change as input, and output the description for the change. This can be viewed as the dual task of program repair, as many code changes and their accompanying commit messages concern bug fixing. Tao et al. (2021) provide a survey on methods and datasets for this task up to 2021.
3.1.4 Code-to-Pattern
Code-to-pattern tasks conduct classification on code. Related methods are listed in Figure 8.
- Type prediction aims to predict the type of dynamic programming languages such as Python and JavaScript. It has been used as a pretraining objective for code language models (Wang et al., 2022d), where it is often simplified as a binary tagging task to predict which tokens in the code are identifiers (Wang et al., 2021c; e).
- Defect detection predicts whether the input code is buggy or not, and is a standard single-sentence classification task. Nong et al. (2023); Steenhoek et al. (2023); Bi et al. (2023); Harzevili et al. (2023) provide surveys on this task.
- Clone detection predicts whether or not two pieces of code are clones of each other. In software engineering there exist four types of code clones, and the most challenging type to identify is semantic clones, i.e. syntactically dissimilar code that have the same functionality. As this task can be viewed as a two-sentence classification task, BERT-style language models have been widely applied to it. Svajlenko & Roy (2020) and Zhang & Sakurai (2021) provide comprehensive reviews on non-deep-learning based methonds for this task.
- Code reasoning is a recently introduced task for evaluating LLMs, and often comes as a subset of general evaluation benchmarks such as MMLU (Hendrycks et al., 2021b). This task requires the model to reason about the code or algorithms, and answer related questions which are written in multiple-choice format or free-form QA format and may range from conceptual understanding to numerical calculation and complexity analysis.
- Code classification aims to predict the functionality of a piece of code within a predefined set of labels. A very similar task is author identification, which predicts the author of the input code. Both tasks are standard single-sentence classification tasks, and traditional machine learning methods have been widely adopted in them (Kalgutkar et al., 2019), while pretrained language models have seen almost no application.
- Machine code detection is another recently introduced task and aims to predict whether the input code is written by human or generated by machines. Detection of AI generated content (AIGC) has undergone thorough studies in general or specialized domains such as education (Mitchell et al., 2023; Liu et al., 2023j), but it is yet to be examined at scale on code.
3.1.5 Text-to-Text
Text-to-text tasks take text as input, and output text. Related methods are listed in Figure 8.
- Document translation is the automatic translation of code-related documents. Since models, datasets, and prompting strategies for machine translation are abundant in NLP (Vaswani et al., 2017; Goyal et al., 2022; He et al., 2023b), we do not go into detail about this task.
- Log analysis aims to analyze the system logs produced by software products, for example parsing logs into structured templates or finding anomalies from raw logs. Zhu et al. (2019) provide a survey on traditional methods for this task up to 2018, and Chen et al. (2021d) give an empirical comparison between neural network based methods. Zhang et al. (2023d) also cover more recent methods for log parsing, while Landauer et al. (2022) survey methods for anomaly detection in logs.
3.1.6 NLP Point-of-View
Among the previously listed tasks, code synthesis, code translation, code repair, deobfuscation, unit test generation, assertion generation, mutant generation, code summarization, code review, identifier prediction, and commit message geneartion are sequence-to-sequence generation tasks. Formally, each instance of these tasks has a source sequence (e.g. a piece of source code) and a target sequence (e.g. its corresponding summarization), and the language model is tasked to maximize the conditional probability given by (3), where can be either a decoder-only model or an encoder-decoder model. In the former case, and are concatenated. In the later case, is processed by the encoder and is processed by the decoder.
Code completion and code infilling are also generation tasks, and correspond exactly to the two pretraining objectives given in (1) and (3), except that for code infilling only one span in the input is masked. Similarly, cloze test is an understanding task in the same form as (2).
Defect detection, clone detection, code classification, and author identification are sequence classification tasks. In these tasks, a set of labels is defined over the input, and each instance is assigned a label (e.g. for defect detection , while for author identification a possible is {Alice, Bob, John, others}). The model is then tasked to maximize
Type prediction is a token classification task, also known as tagging. In this task, each token is assigned a label , with an example being {int, float, string, bool, non-identifier, other}. The model’s objective is to maximize
The last two tasks - code retrieval and code search - also belong to understanding tasks. In these tasks, each source sequence is paired with a positive target sequence and a set of negative targets . The model’s task is to find a similarity metric such that is larger than .
3.2 Evaluation Metrics
Of the tasks mentioned in Section 3.1, the understanding tasks are similar in form to natural language understanding tasks (Wang et al., 2018a; 2019) and evaluated likewise by metrics such as accuracy, F1 and Mean Reciprocal Rank (MRR), while short generation tasks such as identifier prediction is also evaluated by accuracy of exact matches. Code-to-text tasks are evaluated with common metrics for text generation such as BLEU (Papineni et al., 2002),
Evaluation of tasks involving code generation, on the other hand, is more complicated. Most early works evaluate syntactical correctness, i.e. the percentage of generations that can be successfully parsed. Chen et al. (2018) argue against such metrics and suggest reference match instead, which is the percentage of generations that are exactly the same as the references. Ren et al. (2020) propose CodeBLUE, a variant of BLEU that takes code syntax and semantics into account by evaluating the overlap of abstract syntax tree (AST) and data flow.
As code generation models became more capable over the years, however, these metrics based on content-overlap have been found to be inadequate (Rozière et al., 2020; Hendrycks et al., 2021a; Austin et al., 2021), since functionally equivalent snippets of code can differ dramatically in their lexical forms. Consequently, researchers have turned their attention to functional correctness. One popular example of such metrics is pass@, proposed by Kulal et al. (2019) and refined by Chen et al. (2021b), which is an unbiased estimator of the model’s chance in passing all unit tests of a program with any of generated samples. This metric can be generalized to pass (Li et al., 2022g), which limits the number of model submissions to but allows filtering by unit tests given in the input from samples.
3.3 Program Synthesis
We identify four changes in program synthesis over the years: shift of coding paradigms (from example-based to intention-based), generalization in languages (from domain-specific languages to general-purpose languages), simplification of model architectures (from grammar-guided decoders to general-purpose language models), and application of execution-based feedback.
Many of the early methods for program synthesis are exampled-based (Menon et al., 2013), which means they induce programs from input-output examples, often in domain-specific languages (DSLs) such as FlashFill (Devlin et al., 2017a) and Karel444FlashFill is used in Microsoft Excel for string transformation. Karel is a simple programming language for educational purpose. (Bunel et al., 2018), as these languages are usually simple in syntax and structure.
As code generation models became more capable over the years, researchers started to pay attention to program synthesis in general-purpose programming languages as well. Hearthstone (Ling et al., 2016) and CONCODE (Iyer et al., 2018) are two of the early datasets, representing Python and Java respectively. Each example in Hearthstone is the description of a card in the game and its corresponding class implementation, while examples in CONCODE are simply Java methods paired with their natural-language documentation crawled from public GitHub repositories. Synthesizing programs from their corresponding natural language descriptions has since then become a standard practice in program synthesis, and has led to some of the most widely used benchmarks, such as HumanEval (Chen et al., 2021b), which has even been translated into multiple languages (Cassano et al., 2023a; Zheng et al., 2023a; Muennighoff et al., 2023). Some recent benchmarks use general-purpose languages but focus on specific domains, such as data science (Bavishi et al., 2019; Lai et al., 2023) or Jupyter notebooks (Agashe et al., 2019), while several math reasoning benchmarks have also been converted to programming tasks, including MathQA-Python (Amini et al., 2019; Austin et al., 2021) and GSM8K-Python (Cobbe et al., 2021; Chowdhery et al., 2022; Wang et al., 2023d).
Many early works argue that simply treating program synthesis as a text generation task does not utilize the underlying syntax of programming languages, and thus often use syntax-enhanced decoders to inject the target syntax as prior knowledge (Yin & Neubig, 2017). LLMs, however, have demonstrated that pretrained language models are capable of generating syntactically correct programs without loss of generality. Under this setting, researches start to execute the generated programs and provide feedback to the generation model to inject the prior knowledge of code instead. This has recently led to the popularity of interactive coding, which we discuss in more detail in Section 7.1.
3.4 Repository-Level Evaluation
Most evaluation tasks discussed in Section 3.1 are limited to a single file or even a single function, as cross-file code modeling poses challenges that are beyond the capability of most existing language models. Recently, however, position interpolation techniques (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a) have extended the context window of LLMs to hundreds of thousands of tokens, making it possible to contextualize the evaluation of code modeling within entire repositories. Several works (Shrivastava et al., 2023b; Ding et al., 2022b; Zhang et al., 2023b; Shrivastava et al., 2023a) have studied code completion leveraging repository-level context, and Liu et al. (2023i) propose RepoBench to evaluate such systems. More recently, Bairi et al. (2023) investigate the more challenging tasks of repository-level API migration and temporal editing, and Jimenez et al. (2023) introduce a corresponding benchmark, SWE-bench.
4 General Language Models for Code
Since language models scaled to hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2022), many of them have demonstrated non-trivial coding capability, even if they are not specifically designed or trained for code. Pioneered by Codex, researchers have also found continual pretraining on code to significantly benefit language models’ performance on code555While some works refer to this process as “finetuning on code”, it is still self-supervised in nature. Thus we choose to adopt the term “extra/additional/continual pretraining” in this work to avoid confusion with supervised in-task finetuning or instruction finetuning..
4.1 Off-the-Shelf Language Models
Large language models are often pretrained on trillions of tokens following the scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), and such an amount of text data is often a diverse composite with a non-negligible part of code. The Pile (Gao et al., 2021), for example, includes 95GB of code crawled from GitHub out of its 800GB raw dataset, while the multilingual pretraining dataset ROOTS (Laurençon et al., 2022) also contains 163GB of code spanning 13 programming languages in its 1.6TB compound. As two of the largest open-source pretraining datasets, they have supported many language models with coding ability. GPT-J (Wang & Komatsuzaki, 2021), for example, is reported by Chen et al. (2021b) to demonstrate non-trivial performance on HumanEval, while Scao et al. (2022) report similar results for GPT-NeoX (Black et al., 2022) and BLOOM. LLaMA (Touvron et al., 2023a), whose pretraining dataset includes 328GB code from GitHub, achieves 23.7 pass@ performance on HumanEval, and its successor LLaMA 2 (Touvron et al., 2023b), achieves an even higher score of 29.9.
Closed-source models, on the other hand, perform generally better. LaMDA (Thoppilan et al., 2022) and PaLM (Chowdhery et al., 2022), whose pretraining dataset contains 12.5% and 5% code respectively, achieve 14.0 and 26.2 pass@ performance on HumanEval, while GPT-4 (OpenAI, 2023) set a staggering record of 67.0 (and an early version is reported by Bubeck et al. (2023) to be 82) that until recently has remained higher than any specialized models pretrained or instruction-finetuned for code.
| HumanEval (0) | MBPP (3) | |||
| k=1 | k=100 | k=1 | k=80 | |
| GPT-J | 11.6 | 27.7 | ||
| LaMDA | 14.0 | 47.3 | 14.8 | 62.4 |
| PaLM | 26.2 | 76.2 | 36.8 | 75.0 |
| GPT-NeoX | 15.4 | 41.2 | ||
| BLOOM | 15.5 | 55.5 | ||
| LLaMA | 23.7 | 79.3 | 37.7 | 76.8 |
| GPT-4 | 67.0/82 | |||
| LLaMA 2 | 29.9 | 89.0 | 45.0 | 81.5 |
| Phi-1.5 | 41.4 | 43.5 | ||
| Baichuan 2 | 17.1 | 30.2 | ||
| Qwen | 32.3 | 40.8 | ||
| Codex | 28.8 | 72.3 | ||
| PaLM-Coder | 36.0 | 88.4 | 47.0 | 80.8 |
| PaLM 2-S* | 37.6 | 88.4 | 50.0 | 86.8 |
| Code LLaMA | 53.7 | 94.7 | 56.2 | |
| CodeFuse | 74.4 | 61.0 | ||
More recently, the general trend has been to train smaller models with larger datasets, following the revised scaling law (Hoffmann et al., 2022). Baichuan 2 (Yang et al., 2023a), for example, is a 13B model trained on 2.6T tokens, while Qwen (Bai et al., 2023) is a 14B model trained on 3T tokens. They achieve 17.1 and 32.3 pass@1 on HumanEval, respectively. Li et al. (2023h), however, demonstrate that models as small as 1.3B can acquire coding capability that’s comparable to much larger models while also maintaining a reasonable performance on general text processing and even manifesting some emergent abilities (Wei et al., 2022c) such as chain-of-though reasoning (Wei et al., 2022d). Their model, Phi-1.5, is trained on 21B tokens of textbook data generated by ChatGPT, and 100B tokens of filtered web data from Stack Overflow and Refined Web (Penedo et al., 2023), and attains 41.4 pass@ performance on HumanEval. The exact performance of these models are presented in Table 1.
4.2 Language Models with Additional Pretraining on Code
Along with the seminal benchmark HumanEval, Chen et al. (2021b) kick-started the age of LLM for code with Codex, which are GPT-3 checkpoints pretrained on 100B additional code tokens and one of the earliest multi-billion models for code. Following their work, other researchers have also specialized their LLMs on code with additional pretraining. Chowdhery et al. (2022) train PaLM on 7.8B additional code tokens to obtain PaLM-Coder, setting new state-of-the-art on HumanEval and MBPP (Table 1) that are only broken later by its successor PaLM 2-S*, the smallest version of PaLM 2 (Anil et al., 2023) further trained on an undisclosed amount of code. Similarly, Lewkowycz et al. (2022) train PaLM on 38.5B tokens of arXiv papers and mathematical content, while Rozière et al. (2023) train LLaMA 2 (Touvron et al., 2023b) on more than 500B code tokens to acquire Code LLaMA, whose performance on HumanEval surpasses all previous LMs except GPT-4 (Table 1). Liu et al. (2023b) further train Code LLaMA with multi-task finetuning (MFT) to introduce CodeFuse-CodeLLaMA, achieving 74.4 pass@1 on HumanEval and surpassing even the performance of GPT-4 published in OpenAI (2023).
While almost all of these models are Transformer decoders pretrained with CLM, several architectural modifications have been introduced along the way, as we noted in Section 2.5. All these models use pre-norm, and GPT-J introduces parallel attention, which is later adopted by PaLM, GPT-NeoX, and Phi-1.5. PaLM introduces MQA and RoPE into LLMs, and RoPE is now employed by most language models, including GPT-NeoX, two generations of LLaMA, Qwen, and the 7B version of Baichuan 2. BLOOM and the 13B version of Baichuan 2, however, use ALiBi for position embeddings, while LLaMA 2 and Code LLaMA adopt GQA instead of MHA or MQA. In Section 5, we show that specialized models pretrained exclusively on code have also followed these advancements closely.
5 Specialized Language Models for Code
As pretrained Transformers such as GPT and BERT achieved remarkable success in natural language processing, such model architectures, learning paradigms, and training objectives were soon adopted by the software engineering community to produce specialized models for code understanding and generation. In this section, we first review common datasets used for pretraining code language models (Section 5.1), and then dive into the complex family of code LMs by their model architecture: encoder-only models (Section 5.2), encoder-decoder models (Section 5.3), decoder-only models (Section 5.4), UniLM (Section 5.5), and diffusion models (Section 5.6). Lastly, in Section 5.7 we also illustrate the current trend of applying more recent techniques in NLP, such as instruction tuning (Wei et al., 2022b; Sanh et al., 2022; Chung et al., 2022) and reinforcement learning (Ouyang et al., 2022) to code processing. An overview of these pretrained models are provided in Table 3.
5.1 Training Dataset for Code
While text data for pretraining language models are often crawled from the web and must undergo meticulous and often aggressive preprocessing (Raffel et al., 2020), code data come naturally as whole documents from public GitHub repositories. Even better, they come with readily available quality indicators such as the count of stars or forks (although Allal et al. (2023) suggest that star count correlates poorly with downstream performance). As a result, many large-scale code pretraining datasets have been introduced, including CodeSearchNet (Husain et al., 2019), CodeParrot (Tunstall et al., 2022), and the Stack (Kocetkov et al., 2022), totaling 20GB, 50GB and 3TB of code documents respectively (Table 2).
While these datasets are meant for training code models, it should be noted that code is ultimately a special form of natural language, as the vocabulary of most programming languages is a small subset of English. Besides, high-quality code is often interleaved with natural language comments or documentations, which also enables models to acquire certain knowledge of general text representation. In fact, of the 6.5M functions in CodeSearchNet, 2.3M are paired with natural language documentation, allowing models to train explicitly on such bimodal data.
Compared with natural language, another byproduct of scraping code from GitHub is commit histories, which consist of code before commit, code after commit, and a short message describing the commit, which can loosely serve as an instruction for language models. Muennighoff et al. (2023) utilize this feature and construct a 2GB dataset CommitPackFT containing 742K samples of instruction data for code, obviating the need of extensive human labor that’s required to construct natural language instructions (Sanh et al., 2022; Wang et al., 2022f).
Apart from bimodal training and instruction finetuning, another recent trend in constructing code dataset is synthesizing data with powerful models such as ChatGPT. While this method is originally proposed for generating instruction data in natural language (Wang et al., 2023c; Honovich et al., 2023), Gunasekar et al. (2023) take one step further and synthesize 1B tokens of Python textbooks and coding exercises to pretrain a 1.3B model, achieving state-of-the-art results on HumanEval that’s comparable to much larger models trained on significantly larger datasets.
| Dataset | Size (GB) | Files (M) | # PL |
| CodeSearchNet | 20 | 6.5 | 6 |
| The Pile | 95 | 19 | - |
| CodeParrot | 1K | 115 | 30 |
| The Stack | 3136 | 317 | 30 |
| ROOTS | 163 | 15 | 13 |
Date Model Arch. Size Vocab Context PE Preci-sion Atten. Type Parallel Atten. Pre-Norm Flash Atten. Init. from Objectives Dataset Training PL Inst. 2019-12 CuBERT BERT 350M 50K 1024 absolute fp32 MHA - MLM + NSP 9.3B 93B 1 Google 2020-02 CodeBERT RoBERTa 125M 50K 512 absolute fp32 MHA RoBERTa MLM + RTD 20GB 105B 6 Microsoft 2020-09 GraphCode-BERT RoBERTa 125M 50K 640 absolute fp32 MHA CodeBERT MLM + Edge Prediction + Node Alignment 20GB 131B 6 Microsoft 2021-08 SynCoBERT RoBERTa 125M 50K 512 absolute fp32 MHA CodeBERT MLM + IP + AST Edge Prediction + CL 20GB 7B 6 Huawei 2021-10 DISCO BERT 100M 20K 512 absolute fp32 MHA - MLM + Node Type MLM + CL 1.8GB 2 Columbia & IBM 2022-05 Code-MVP RoBERTa 125M 50K 512 absolute fp32 MHA GraphCode-BERT MLM + Type Inference + CL 2GB 39B 1 Huawei 2022-10 SCodeR RoBERTa 125M 51K 1024 absolute fp32 MHA UniXcoder CL 20GB 6 Microsoft 2020-05 GPT-C GPT-2 374M 60K 1024 absolute MHA ✓ - CLM 11B 270B 4 Microsoft 2021-02 CodeGPT GPT-2 124M 50K 1024 absolute fp32 MHA ✓ GPT-2 CLM 2GB 1 Microsoft 2022-02 PolyCoder GPT-NeoX 160M-2.7B 50K 2048 RoPE fp32 MHA ✓ - CLM 254GB 39B 12 CMU 2022-03 CodeGen-Multi(Mono) GPT-J 350M-16.1B 50K 2048 RoPE fp16 MHA ✓ ✓ - CLM 1.6TB(1.8TB)/ 506B(577B) 1T(1.2T) 6(1) Salesforce 2022-04 InCoder GPT-3 6.7B 50K 2048 Cosine fp32 MHA ✓ - Causal Masking 204GB 52B 28 Meta 2022-06 PyCodeGPT GPT-Neo 110M 32K 1024 absolute fp32 MHA ✓ - CLM 96GB 100B 1 Microsoft 2022-07 PanGu-Coder PanGu- 317M-2.6B 42K 1024 absolute MHA ✓ - CLM 147GB 230B 1 Huawei 2023-01 SantaCoder GPT-2 1.1B 49K 2048 absolute fp32 MQA ✓ - FIM 268GB 236B 3 BigCode 2023-03 CodeGeeX PanGu- 13B 52K 2048 absolute fp16 MHA ✓ - CLM 158B 850B 23 Tsinghua 2023-05 StarCoder GPT-2 15.5B 49K 8192 absolute fp32 MQA ✓ ✓ - FIM 815GB 1T 86 BigCode 2023-05 Jam GPT-2 350M 50K 256 absolute bf16 MHA ✓ - CLM 36GB/20B 9B 1 U. Notre Dame 2023-06 Phi-1 GPT-J 1.3B 51K 2048 RoPE fp16 MHA ✓ ✓ ✓ - CLM 7B 53B 1 Microsoft 2023-10 CodeFuse GPT-J 350M-13B 101K 4096 RoPE fp16 MHA ✓ ✓ ✓ - CLM 1.6TB / 1T 40+ Ant Group 2023-10 CodeShell GPT-2 7B 70K 8192 RoPE bf16 GQA ✓ - CLM 500B Peking U. 2020-10 PyMT5 GPT-2 374M 50K 1024+1024 absolute fp16 MHA ✓ - SC 27GB 1 Microsoft 2021-02 Mastropaolo et al. T5 60M 32k 512+512 T5 bf16 MHA ✓ - SC 1GB 1 Universita della Svizzera italiana 2021-02 DOBF 250M 50K 512+512 absolute fp16 MHA - MLM + Deobfuscation 45GB 2 Meta 2021-03 PLBART BART 140M 50K 1024+1024 absolute fp32 MHA - DAE 655GB / 71B 210B 2 UCLA & Columbia 2021-09 CodeT5 T5 60M-220M 32K 512+256 T5 fp16 MHA ✓ - SC + IP + Masked IP + Text2Code + Code2Text 25GB 8 Salesforce 2022-01 SPT-Code BART 262M 80K 512+512 absolute fp16 MHA - NSP + SC + Method Name Prediction 20GB 6 Nanjing U. 2022-02 AlphaCode 300M-41B 8K 1536+768 bf16 MQA - MLM + CLM 715GB 967B 13 DeepMind 2022-06 NatGen T5 220M 32K 512+256 T5 fp16 MHA ✓ CodeT5 Naturalization 26GB 14B 8 Columbia & UC Davis 2022-12 ERNIE-Code mT5 580M 250K 1024+1024 T5 bf16 MHA ✓ mT5 SC + Text2Code + Code2Text + Machine Translation 197B 6 Baidu 2023-05 CodeT5+ T5/GPT-3 220M-16B 50K 2048+2048 absolute fp16 MHA ✓ ✓ CodeGen-mono SC + CLM + CL + Text2Code + Code2Text 52B 9 Salesforce 2020-12 CugLM BERT 51M 50K 128 absolute fp32 MHA - MLM + NSP + CLM 8M 1.2B 2 Peking U. 2022-03 UniXcoder RoBERTa 125M 51K 1024 absolute fp32 MHA - MLM + CLM + SC + CL + Code2Text 20GB+ 839B 6 Microsoft
5.2 Encoders
Pretrained Transformer encoders such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019d), and ELECTRA (Clark et al., 2020) have attained impressive results on natural language understanding tasks, and these methods were soon introduced into code processing after their advent. Kanade et al. (2020) replicate the training procedure of BERT on a code corpus to produce CuBERT, showcasing its superior performance over LSTM (Hochreiter & Schmidhuber, 1997) and non-pretrained Transformers. Feng et al. (2020), on the other hand, train CodeBERT with MLM and ELECTRA’s RTD on CodeSearchNet. They also utilize the explicit text-code pairs in CodeSearchNet, and use them respectively as the first and second segment in BERT’s input. When using CodeBERT to initialize the encoder part of a vanilla Transformer for sequence-to-sequence generation tasks such as code summarization, they observe a moderate performance gain over non-pretrained baselines.
Apart from these standard training objectives, many auxiliary objectives specifically designed for code have also been introduced. GraphCodeBERT (Guo et al., 2021a) and SynCoBERT (Wang et al., 2021c) both extract graphs from the source code (data flow graph and abstract syntax tree, respectively) and train the models to predict the typological relations between the nodes, while SynCoBERT and Code-MVP (Wang et al., 2022d) also add type inference to their pretraining stage in the form of tagging. Another common objective is contrastive learning: SynCoBERT and Code-MVP contrast between different views of the input (such as code, comment, AST, and transformed code), while DISCO (Ding et al., 2022a) constructs positive sample pairs by semantic-preserving transformations such as obfuscation, and negative pairs by injecting artificial bugs.
5.3 Encoder-Decoders
In NLP, pretrained Transformer encoder-decoders such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020) have also left a notable mark in the past few years’ advancement in language modeling. T5, for example, unifies all textual tasks into a sequence to sequence format and sets new records on GLUE (Wang et al., 2018a) and SuperGLUE (Wang et al., 2019). Compared with encoder-only models, encoder-decoders are naturally more powerful as they can be used for conditional text generation, while their encoder part can always be taken out to perform tasks that require an encoder-only architecture, such as regression (Tay et al., 2023).
Inspired by these advantages of encoder-decoder architecture, many such models have been proposed for code processing. PyMT5 (Clement et al., 2020) and Mastropaolo et al. (2021) replicate the pretraining and multi-task finetuning process of T5 on code corpus, while Ahmad et al. (2021) introduce PLBART, a BART pretrained on 655GB combined data of Java, Pyhton, and natural language. Lachaux et al. (2021) argue that MLM could be too easy a task for programming languages as identifier names often occur multiple times in a single context window, and propose a deobfuscation pretraining objective, where the model is trained to convert obfuscated code back to its original form. Related to this method, we note that meaningful variable names have also been found to have a positive impact on the code generation process of large language models (Chen et al., 2022).
Building on these early works, Wang et al. (2021e) propose CodeT5, which is pretrained alternatively with 1) T5’s original span corruption; 2) identifier tagging (where each token in the code input is tagged as either identifier or non-identifier); 3) masked identifier prediction (a special form of span corruption where all identifiers are masked); and 4) text-to-code & code-to-text generation. Its successor, CodeT5+ (Wang et al., 2023d), take inspiration from UL2 (Tay et al., 2023) and introduce causal language modeling (CLM) into pretraining, along with additional contrastive objectives based on text-code matching.
AlphaCode (Li et al., 2022g) is also trained with multiple objectives, where the encoder is trained with MLM and the decoder is trained with CLM, with architecture modifications such as shallow-encoder & deep-decoder, multi-query attention (Shazeer, 2019), and being much larger than CodeT5 (up to 41B parameters). NatGen (Chakraborty et al., 2022a), on the other hand, is pretrained with a "naturalization" objective similar to deobfuscation: semantically equivalent but unnatural code is generated by predefined operations such as loop transformation, dead code injection, and variable renaming, and the model is pretrained to translate these unnatural code back to its original form. We note that some of these models are built on previous works. For example, NatGen is initialized with CodeT5, while the largest version of CodeT5+ is initialized from a decoder-only model, CodeGen (Nijkamp et al., 2023b).
Apart from these general pretraining objectives, several works have also trained Transformer encoder-decoders with a focus on code translation, which is a natural application of Transformer models in code as the Transformer architecture was originally proposed by Vaswani et al. (2017) for machine translation (MT). However, unlike natural languages, where parallel corpus across two or more human languages exist in abundance, there is little parallel data for code. To tackle this issue, Rozière et al. (2020) propose Transcoder, which first pretrains an encoder with XLM (Conneau & Lample, 2019), and then initializes a vanilla Transformer with this encoder and continue to pretrain it with Denoising Auto-Encoding (DAE, Lewis et al., 2020) and back translation (Sennrich et al., 2016), while its follow-up work (Szafraniec et al., 2023) also utilize language-independent intermediate representations to enhance this process, which we discuss in more detail in Section 6.
Apart from training data and objectives, these models mostly keep to the original architectures proposed by the NLP community, as shown in Table 3. Models based on BART, for example, use post-normalization and learnable absolute position embeddings, while those based on T5 use its simplified relative position embeddings and pre-normalization.
5.4 Decoders
After the monumental debut of GPT-3 (Brown et al., 2020) and the discovery of in-context learning, decoder-only Transformer models have become dominant in language modeling (Rae et al., 2021; Hoffmann et al., 2022; Chowdhery et al., 2022; Scao et al., 2022; Touvron et al., 2023a; b, inter alia). Many models similarly pretrained with CLM have also emerged in code processing, such as GPT-C (Svyatkovskiy et al., 2020), CodeGPT (Lu et al., 2021), PolyCoder (Xu et al., 2022), CodeGen (Nijkamp et al., 2023b), PyCodeGPT (Zan et al., 2022), Pangu-Coder (Christopoulou et al., 2022), CodeGeeX (Zheng et al., 2023a), Jam (Su et al., 2023), Phi-1 (Gunasekar et al., 2023), CodeFuse (Di et al., 2023), CodeShell666https://github.com/WisdomShell/codeshell, and DeepSeek Coder777https://github.com/deepseek-ai/DeepSeek-Coder. Of these models, several alternative training objectives have been experimented with, such as MLM and Masked CLM888In their paper, MLM is conducted by replacing tokens in the input with <mask> and predicting it from only the left context, while Masked CLM is performed by adding a <mask> in the input and predict the the next token from it. Both tasks do not change the attention mask patterns of the model. in Pangu-Coder, but are found to underperform compared with CLM-only training. Zan et al. (2022) also propose continual training on sketches, where the model learns to first generate a sketch of a program and then the actual code. Notably, Gunasekar et al. (2023) present Phi-1, a 1.3B small model trained on a dataset of only 7B tokens consisting of 6B tokens from StackOverflow and 1B synthetic data generated by ChatGPT but achieving 50.6 pass@ on HumanEval and 55.5 pass@ on MBPP, comparable to much larger (both in model size and training data size) models such as Code LLaMA or PaLM 2.
| Model | Size | HumanEval | MBPP |
| PolyCoder | 2.7B | 5.6 | - |
| CodeGen-Mono | 16.1B | 29.3 | 35.3 |
| InCoder | 6.7B | 15.2 | 19.4 |
| PyCodeGPT | 110M | 8.3 | - |
| Pangu-Coder | 2.6B | 23.8 | 23.0 |
| SantaCoder | 1.1B | 14.0 | 35.0 |
| CodeGeeX | 13B | 22.9 | 24.4 |
| StarCoder | 15.5B | 33.6 | 52.7 |
| CodeT5+ | 16B | 30.9 | - |
| Phi-1 | 1.3B | 50.6 | 55.5 |
| CodeFuse | 13B | 24.8 | - |
| InstructCodeT5+ | 16B | 35.0 | - |
| WizardCoder | 15.5B | 57.3 | 51.8 |
| Pangu-Coder 2 | 15.5B | 61.6 | - |
| OctoCoder | 15.5B | 46.2 | - |
| CodeFuse-SFT | 13B | 37.1 | - |
| GPT-4 | - | 67.0/82 | - |
| PaLM 2-S* | - | 37.6 | 50.0 |
| Code LLaMA | 34B | 53.7 | 56.2 |
| Phi-1.5 | 1.3B | 41.4 | 43.5 |
Although Christopoulou et al. (2022) report denoising objectives to underperform in decoder-only models, there have been other works that successfully combine denoising or multi-task pretraining with decoder architecture. Incoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023g) are all trained with fill-in-the-middle (FIM) objective, also referred to as causal masking by Fried et al. (2023), which is essentially span corruption (Raffel et al., 2020) adopted to decoder-only architecture. One of the visible advantages of these infilling objectives is that they inject the models with the ability to fill in blanks in the middle of input code at inference time, while CLM allows only for autoregressive generation. As Table 4 shows, however, these objectives also lead to higher performance on downstream tasks when compared with CLM-only models such as CodeGen, although the exact benefits of infilling training remains controversial (Nijkamp et al., 2023a).
Observing Table 3, it is clear that decoder-only models for code have generally followed the practices in NLP more closely, when compared with other model architectures. All these models use pre-normalization, while MQA, RoPE, and parallel attention have also been adopted by several models. Notably, the three most recent models - StarCoder, Phi-1, and CodeFuse - also employ FlashAttention to improve model throughput.
5.5 UniLMs
Following UniLM (Dong et al., 2019) in NLP, several works in code processing have also pretrained this fourth family of Transformer models on code. CugLM (Liu et al., 2020) is trained with both CLM and MLM + NSP via alternating attention masks, while UniXcoder is trained with CLM, MLM, Span Corruption (in Prefix LM style) along with auxiliary objectives including contrastive learning and text-code mutual generation. Both two models, however, are relatively small in size, and whether or not this architecture is suitable for code processing is yet to be explored at scale.
5.6 Diffusion Models
Currently the Transformer architecture dominate text generation, but several works (Li et al., 2022c; Lin et al., 2023) have also adopted Diffusion Models (Ho et al., 2020) from computer vision for text generation. Recently CodeFusion (Singh et al., 2023) also introduces diffusion models into code modeling, and demonstrates that a 75M diffusion model can outperform StarCoder, CodeT5+, and GPT-3 on three code synthesis datasets.
5.7 Instruction Finetuning and Reinforcement Learning for Code
In natural language processing, training models on a diverse set of tasks with instruction prefix, known as instruction finetuning, has been shown to unlock the ability of cross-task generalization (Ouyang et al., 2022; Chung et al., 2022; Iyer et al., 2022). At first, these instruction data samples are manually compiled or crowd-sourced (Wei et al., 2022b; Sanh et al., 2022), but later researches find LLM-generated instructions to be sufficient (Wang et al., 2023c; Honovich et al., 2023).
Following these works in natural language, researchers from the code community have applied instruction tuning to their models as well. Wang et al. (2023d) finetune CodeT5+ with 20K instruction data generated by InstructGPT (Ouyang et al., 2022) to obtain InstructCodeT5+. WizardCoder (Luo et al., 2023) follows the methods of WizardLM (Xu et al., 2023a) to evolve 20K code Alpaca (Taori et al., 2023) samples into a 78K dataset and uses it to finetune StarCoder. Pangu-Coder 2 (Shen et al., 2023) also uses WizardLM’s Evol-Instruct to generate 68K instruction samples from 20K code Alpaca, but also introduces reinforcement learning via Rank Responses to align Test & Teacher Feedback (RRTF). OctoCoder (Muennighoff et al., 2023), on the other hand, takes a different path and uses Git commit histories as instruction data to finetune StarCoder and CodeGeeX2. More recently, CodeFuse (Di et al., 2023) also employs multitask-finetuning and explicitly introduces multiple downstream tasks into their instruction data. The performance of these instruction finetuned code models can also be found in Table 4.
In NLP, another technology closely related to instruction finetuning is reinforcement learning from human feedback (RLHF), which has played a significant role in aligning LLMs with human values (Ouyang et al., 2022; Bai et al., 2022). The merit of reinforcement learning is that it can incorporate non-differentiable reward signals into training, such as BLEU (Bahdanau et al., 2017) and human preference (Christiano et al., 2017), but the human feedback required in aligning LLMs often involves extensive labor on annotation. In comparison, applying reinforcement learning to code models has a natural advantage, as compilers can be used for automatically generating feedback for code samples produced by language models.
CodeRL (Le et al., 2022) is one such model, which defines four levels of rewards for each generated program (viz. compile error, runtime error, unit test failure, pass) as well as fine-grained token-level reward estimated by a critic model. The actor model, which is an extention of CodeT5, is then trained with REINFORCE algorithm (Williams, 1992). Similarly, CompCoder (Wang et al., 2022c) and PPOCoder (Shojaee et al., 2023) train CodeGPT and CodeT5 respectively with proximal policy optimization (Schulman et al., 2017), while RLTF (Liu et al., 2023d) proposes fine-grained feedback based on the error information and location provided by the compiler, as well as adaptive feedback that takes the ratio of passed test cases into account.
6 Code Features for Language Models
A major difference between programming languages and natural languages is that the former is artificially defined to be precise and unambiguous, and need to be compiled (or interpreted) without error before execution. This allows for a much larger flexibility in designing pretraining objectives on code, beside lexical manipulations such as CLM, MLM, and Span Corruption. A similar trend can be observed in the last years before neural networks were introduced into mainstream NLP literature (Sutskever et al., 2014; Bahdanau et al., 2015), when researchers in the MT community utilized alternative views of text such as syntactic features to improve the performance of SMT systems (Galley et al., 2006; Chiang, 2007). These features, however, are not universally applicable or even agreed upon, and often result in highly complicated systems (for example, the size of English part-of-speech tagging’s label set may range from dozens to hundreds).
Programming languages, however, fare much better in these aspects. Each mainstream programming language, such as C, Python, and Java, comes with readily available compiler toolkits that allow for easy and accurate extraction of semantic information such as Abstract Syntax Tree (AST), language-independent Intermediate Representation (IR), and auxiliary information such as type of each token and control/data flow graph (CFG/DFG). Thus, in the context of Transformer-based language modeling for code, many works have incorporated these features into their training procedure.
6.1 Abstract Syntax Tree and Intermediate Representation
AST is one of the most common intermediate results of the compiling process, where a program is parsed into a tree of operations and their operands. Before the popularization of Transformer in the code processing community, there had been works such as InferCode (Bui et al., 2021a) that processes these representations with special network architectures like Tree-Based CNN and conducts self-supervised pretraining by predicting subtrees.
TreeBERT (Jiang et al., 2021b) is one of the first attempts to take AST into the Transformer-based pretraining-finetuning framework. It’s a Transformer encoder-decoder pretrained with Tree MLM and Node Order Prediction, where the encoder takes a set of constituent paths in the AST as input (with each token being a path, which is the concatenation of its nodes’ representations) while the decoder takes the code as input. Tree MLM is then performed by masking certain nodes in a path representation and its corresponding code tokens in the decoder input, while Node Order Prediction is accomplished by swapping nodes in a path and predicting it with a [CLS] token similar to BERT.
The method used by TreeBERT, however, is complicated and does not scale well. Later works mostly opt to first process AST into a text sequence and treat it like a normal part of the input. Wang et al. (2021c), for example, process AST with depth-first traversal and concatenate it with code and comment, and then train SynCoBERT (which, unlike TreeBERT, is actually a BERT-like encoder-only model) with four objectives: 1) MLM; 2) identifier tagging; 3) AST edge prediction (predicting whether there exists an edge between two AST nodes from the dot product of these nodes’ representations); and 4) contrastive learning over i) code and AST pairs, as well as ii) text and code-AST pairs. Similarly, SPT-Code (Niu et al., 2022), a Transformer encoder-decoder, takes the concatenation of code, sequentialized AST, and text as input, and is pretrained with 1) span corruption; 2) code-AST prediction (NSP with one segment being code and one segment being AST); and 3) method name generation, a special form of span corruption where a method name is masked. Different from other works, however, they do not take the docstrings as the text segment in their input, but instead concatenate all method names appearing in the code as a succinct natural language description. Likewise, UniXcoder (Guo et al., 2022) takes flattened AST instead of source code as its input during training.
In the compiling pipeline, AST is usually followed by language-independent intermediate representations, such as LLVM IR (Lattner & Adve, 2004). Such features’ independence from specific programming languages makes them suitable candidates for translation pivots, as is English in machine translation of low-resource natural languages (Leng et al., 2019). Szafraniec et al. (2023) take advantage of this characteristic and extend Transcoder (Rozière et al., 2020) with translation language modeling (Conneau & Lample, 2019) over code and IR, as well as IR generation from code. They also investigate other objectives such as IR decompilation (i.e. generating code from IR) and IR pivot (i.e. directly generating code in one language from the IR of another language), both showing promising results.
6.2 Control Flow and Data Flow
While AST and IR have proved to be useful information in certain tasks such as code translation, they are static by nature, just like the source code, and may fail to capture semantic properties of code that are only revealed at runtime (Wang & Su, 2020). Such semantics, however, are contained in dynamic features such as control flow and data flow. Similar to AST, specialized networks were used to process such information before the rise of pretrained Transformers, such as Message Passing Neural Network used by ProGraML (Cummins et al., 2021). Unlike AST, however, even after pretrained Transformers became dominant few works have looked in this direction.
GraphCodeBERT (Guo et al., 2021a) is one of such works, which creates special tokens and position embeddings for variables in the flow graph, and concatenates the variable sequence after text and source code to construct model input, with tailored attention masks on the code and variable segments: tokens from code segment and variable segment can attend to each other if and only if the variable is identified from the code token, and for tokens within the variable segment, is allowed to attend to if there is a direct edge from to in the dataflow. The model is then pretrained with MLM in combination with edge prediction and node alignment, both of which are accomplished by binary classification from the dot product of two tokens’ representations (one from code segment and one from variable segment for node alignment, and both from variable segment for edge prediction).
6.3 Type
Apart from AST, IR, and data flow, type information has also been used to aid language models in processing code. CugLM (Liu et al., 2020), for example, uses type information during finetuning to aid in the prediction of tokens for unidirectional MLM (i.e. MLM with unidirectional attention mask): the type of a masked token is first predicted from the final Transformer layer’s representation, and then the token itself is predicted based on both the hidden representation and predicted type. In contrast, both CodeT5 (Wang et al., 2021e) and SynCoBERT (Wang et al., 2021c) include identifier tagging in their pretraining objectives, which can be viewed as coarse-grained type prediction.
6.4 Program Transformation
As we have shown in Section 5.3, function-preserving program transformations have proven to be important techniques in pretraining code language models. Obfuscation is one instance of program transformation, and others include loop transformation (for-while), condition transformation (if-switch), dead code injection (e.g. if True: pass), and statement swapping. DOBF (Lachaux et al., 2021) and NatGen (Chakraborty et al., 2022a) are two code language models pretrained to recover the original program from such transformations, while Wang et al. (2022a) also apply program transformations during the finetuning stage of language models to make them more robust to transformed test samples.
Notably, Wang et al. (2022d) integrate many of the aforementioned features into Code-MVP: source code, docstrings, AST, CFG, and transformed source code via identifier renaming, loop exchange, and dead code insertion. The model, initialized from GraphCodeBERT, is then trained with MLM, fine-grained type prediction, and contrastive learning across different views, such as text vs. code, code vs. AST, and code vs. CFG.
7 LLMs in Software Development
As language models set new records on software engineering benchmarks, software engineering technologies are also expanding the boundaries of language models in return, and have subsequently led them into real-world development cycles.
7.1 LLMs Extended with Coding Tools
Researches in the NLP community have shown that LLMs can learn to use external tools such as calculators, MT systems, and search engines (Thoppilan et al., 2022; Schick et al., 2023). As such, interpreter has been used to augment LLMs in complex reasoning tasks. PAL (Gao et al., 2023b) and PoT (Chen et al., 2022) both extend Codex with Python interpreters for numerical calculations, while ViperGPT (Surís et al., 2023) extends it further by calling vision APIs to extract information from visual input and answer related questions.
Apart from alleviating the burden of numerical calculation in abstract reasoning tasks, interpreter also provides feedback on the process of code generation itself, together with unit tests. CodeT (Arteca et al., 2022) and TiCoder (Bareiß et al., 2022) use Codex to generate unit tests, which are run against generated code samples to improve the model’s performance on code synthesis. Similarly, TransCoder-ST (Rozière et al., 2022) augments TransCoder and DOBF with external unit tests for code translation, while InterCode (Yang et al., 2023d) explores interactive coding with execution feedback. In Section 5.7, we have also shown that the execution results on unit tests serve as natural supervision signals for reinforcement learning on code.
Notably, in March 2023 OpenAI also released an interpreter plugin for ChatGPT999https://openai.com/blog/chatgpt-plugins#code-interpreter, which can accept file inputs from users, generate code according to user instructions, and provide feedback via real-time execution. Zhou et al. (2023a) show that this feature allows GPT-4 to self-debug.
A topic closely related to tool using in LLM researches is planning as intelligent agents, which has been shown to enhance LLMs’ capability both theoretically and empirically (Feng et al., 2023). Ruan et al. (2023) find that LLMs can plan to solve complex tasks using external SQL generators and Python generators, while CodePlan (Bairi et al., 2023) demonstrates they can perform repository-level coding via adaptive planning.
Another stream of works use LLMs to create multi-agent systems for code generation, such as self-collaboration (Dong et al., 2023), ChatDev (Qian et al., 2023), and MetaGPT (Hong et al., 2023). In these frameworks, multiple LLMs are prompted to play distinct roles such as programmer, reviewer, and manager. These roles interact with each other, breakdown code generation into different phases (e.g. designing, coding, testing, and documenting), and collaborate to complete complex tasks.
7.2 LLMs Integrated into Software Development
With the increase in LLMs’ interactive coding capability, researchers have also started to integrate them into each and every process of software development.
Auto code completion is one of the earliest applications of language models in software development, as they require only the ability to predict the next token. Even before language models scaled to billions of parameters, there had been integration of completion systems such as Pythia (Svyatkovskiy et al., 2019) and IntelliCode (Svyatkovskiy et al., 2020) into popular IDEs.
Recently, however, the application of code language models have transcended simple code completion. GitHub Copilot is arguably one of the most popular AI code assistants, with diverse features including code generation, vulnerability detection, and license management101010https://github.com/features/copilot, while CodeFuse (Di et al., 2023) also integrates code generation, code translation, code commenting, and testcase generation into a single IDE extension. As code language models become larger, however, their client-side deployment and real-time performance also raise new challenges.
As LLMs continue to advance, building applications on top of them is also evolving into a consequential task itself. Many open-source frameworks for such applications have been released, including LangChain111111https://www.langchain.com/, AutoGPT121212https://github.com/Significant-Gravitas/AutoGPT, and WorkGPT131313https://github.com/team-openpm/workgpt. These frameworks provide abstractions over language models for developers, and are actively revolutionizing the entire process of software development even as this survey is being finalized.
8 Conclusion and Challenges
In this work, we systematically reviewed the history of code processing with pretrained Transformer language models, and highlighted their relations and comparisons to models pretrained on general domains. The advancement in code modeling generally follows the history course of NLP, evolving from SMT models, to NMT models, and then to finetuning pretrained Transformers and lastly to few-shot application of LLMs and even autonomous agents in real-world production. Unlike natural languages, the nature of code makes it easy to extract auxiliary information from alternative views, and to utilize interpreter and unit tests for automatic feedback.
With these in mind, we identify several challenges in the current development of code modeling.
- Comprehensive benchmarks to push code LLMs to the next stage. The widely used HumanEval benchmark plays a key role in the evolution of Code LLMs. However, it is relatively small and its scoreboard has been manipulated to near perfect, which does not exactly reflect real-world behaviors. Many other benchmarks for Code LLMs have been proposed, but they are still not comprehensive enough to reflect production-level requirements. The community is eager for a new standard benchmark after HumanEval to further boost the progress of Code LLMs to the next stage.
- Acquisition of high-quality data. With Gunasekar et al. (2023) achieving SOTA performance with a 1.3B model trained on textbook data, we believe the selection of training data and utilization of synthetic data will be ever more prominent in the near future, for both self-supervised pretraining and supervised finetuning.
- Integration of code features into language models. As we noted in Section 6.2, CFG and DFG are yet to be employed at scale in code language modeling. The few works that do employ data flow make changes to the models’ attention masks, which severely limits their cross-task generalization and scaling ability. We believe the seamless integration of such features into textual input is worth researching in the future.
- Application of LLMs in more code downstream tasks. As we have pointed out in Section 3, current evaluation of LLMs’ coding capability is focused on program synthesis, and Figure 6 clearly shows that tasks related to software testing (viz. unit test generation, assertion generation, mutant generation, and fuzzing) and deobfuscation have seen few application of LLMs. Besides, since the context window of LLMs are currently quite limited, generation tasks such as program synthesis and code translation are yet to be applied beyond method level. In Section 3.4, we have listed several works on repository-level code completion and temporal editing, and we believe the application of LLMs in more repository-level tasks will become a hot research topic in the future.
- Alternative model architectures and training objectives. In Table 3, we have shown that many code language models are pretrained with auxiliary objectives specific to code, but these models all belong to the encoder-only or encoder-decoder family, while decoder-only models are yet to be augmented with alternative objectives. Also, as pioneered by Singh et al. (2023), we believe diffusion models will find its ground in code modeling in the future.
- Building code LLM ecosystem for full-life-cycle of software development. While the academia have witnessed an abundance of code models, most have been deployed in the coding stage as IDE plugins while neglecting other stages in the life-cycle of software development. In Section 7.2 we mentioned several inspiring examples, and we are hoping to see more applications of code LMs throughout the full life-cycle of software development, from requirement analysis to DevOps, eventually leading to full-scale ecosystems like those around PyTorch (Paszke et al., 2019) and Hugging Face141414https://huggingface.co/.
- Safety and ethics issues related to code LLMs. As language models grow in might, they also raise safety concerns including but not limited to data contamination, toxic or biased generation, personal information leak, and hallucinations. In software development, these models should be deployed with extra caution, as their generated code may contain security risks leading to catastrophic results. Pretraining data is also becoming a sensitive topic of ethics, and Kocetkov et al. (2022) take a meaningful step towards this issue by allowing developers to remove their code from the Stack. As synthetic training data becomes widespread, researchers should also proceed with caution about such practice, as the consequence of training AI models with AI generated data is yet to be investigated at scale.
With the presentation of this survey, we hope to provide a global view of language models’ application in software engineering and connect the researches from the two communities. We believe the current surge of LLMs will be ultimately transformed into real world applications, and lead humanity into a brighter future.
Appendix A Benchmarks for Downstrem Tasks
Table 5, 6, 7, 8, 9 list benchmark datasets for code downstream tasks. For the size of these datasets, we use self-reported number whenever available.
Task Date Benchmark Source Size Language Text-to-SQL 1990 ATIS Hemphill et al. (1990); Dahl et al. (1994) 11508 1996 GeoQuery Zelle & Mooney (1996) 877 2000 Restaurants Tang & Mooney (2000) 378 2014-09 MAS Li & Jagadish (2014) 196 2017-02 Yelp Yaghmazadeh et al. (2017) 128 2017-02 IMDb Yaghmazadeh et al. (2017) 131 2017-04 Scholar Iyer et al. (2017) 816 2017-08 WikiSQL Zhong et al. (2017) 80654 2018-06 Advising Finegan-Dollak et al. (2018) 4570 2018-09 Spider Yu et al. (2018c) 10181 2019-06 SParC Yu et al. (2019c) 12726 2019-07 MIMICSQL Wang et al. (2020b) 10000 2019-09 CoSQL Yu et al. (2019b) 15598 2020-05 Criteria-to-SQL Yu et al. (2020) 2003 2020-10 Squall Shi et al. (2020) 11276 2020-10 Spider-Realistic Deng et al. (2021) 508 2021-06 Spider-Syn Gan et al. (2021a) 8034 2021-06 SEDE Hazoom et al. (2021) 12023 2021-06 KaggleDBQA Lee et al. (2021) 400 2021-09 Spider-DK Gan et al. (2021b) 535 2022-05 Spider-SS Gan et al. (2022) 8034 2022-05 Spider-CG Gan et al. (2022) 45599 2023-05 BIRD Li et al. (2023f) 12751 Program Synthesis 2018-02 NL2Bash Lin et al. (2018b) 9305 Bash 2018-08 CONCODE Iyer et al. (2018) 104K Java 2019-10 JuICe Agashe et al. (2019) 1.5M/3725 Python 2021-05 APPS Hendrycks et al. (2021a) 10000 Python 2021-07 HumanEval Chen et al. (2021b) 164 Python 2021-08 MBPP Austin et al. (2021) 974 Python 2021-08 MathQA-Python Austin et al. (2021) 23914 Python 2021-08 PlotCoder Chen et al. (2021c) 40797 Python 2022-01 DSP Chandel et al. (2022) 1119 Python 2022-02 CodeContests Li et al. (2022g) 13610 C++, Python, Java 2022-03 MCoNaLa Wang et al. (2023e) 896 Python 2022-06 AixBench Hao et al. (2022) 336 Java 2022-10 MBXP Athiwaratkun et al. (2023) 12.4K 13 2022-10 Multilingual HumanEval Athiwaratkun et al. (2023) 1.9K 12 2022-10 MathQA-X Athiwaratkun et al. (2023) 5.6K Python, Java, JS 2022-11 ExeDS Huang et al. (2022b) 534 Python 2022-11 DS-1000 Lai et al. (2023) 1000 Python 2022-12 ODEX Wang et al. (2022g) 945 Python 2023-02 CoderEval Yu et al. (2023a) 460 Python, Java 2023-03 xCodeEval Khan et al. (2023) 5.5M 11 2023-03 HumanEval-X Zheng et al. (2023a) 820 Python, C++, Java, JS, Go 2023-06 StudentEval Babe et al. (2023) 1749 Python 2023-06 DotPrompts Agrawal et al. (2023) 10538 Java 2023-09 CodeApex Fu et al. (2023) 476 C++ 2023-09 VerilogEval Liu et al. (2023g) 8645/156 Verilog
Task Date Benchmark Source Size Language Code Translation 2020-06 GeeksforGeeks Rozière et al. (2020) 1.4K C++, Java, Python 2021-02 CodeTrans Lu et al. (2021) 11.8K Java, C# 2021-08 Avatar Ahmad et al. (2023) 9515 Java, Python 2022-06 CoST Zhu et al. (2022b) 132K C++, Java, Python, C#, JS, PHP, C 2022-06 XLCoST Zhu et al. (2022a) 567K C++, Java, Python, C#, JS, PHP, C 2023-03 xCodeEval Khan et al. (2023) 5.6M 11 2023-03 HumanEval-X Zheng et al. (2023a) 1640 Python, C++, Java, JS, Go 2023-08 G-TransEval Jiao et al. (2023) 4000 C++, Java, C#, JS, Python 2023-10 CodeTransOcean Yan et al. (2023) 270.5K 45 Program Repair 2014-07 Defects4J Just et al. (2014) 357 Java 2015-12 ManyBugs Goues et al. (2015) 185 C 2015-12 IntroClass Goues et al. (2015) 998 C 2016-11 BugAID Hanam et al. (2016) 105K JS 2017-02 DeepFix Gupta et al. (2017) 6971 C 2017-05 Codeflaws Tan et al. (2017) 3902 C 2017-10 QuixBugs Lin et al. (2017) 80 Java, Python 2018-05 Bugs.jar Saha et al. (2018) 1158 Java 2018-12 BFP Chakraborty et al. (2022b) 124K Java 2019-01 Bears Madeiral et al. (2019) 251 Java 2019-01 unnamed Tufano et al. (2019a) 21.8K Java 2019-04 BugsJS Gyimesi et al. (2019) 453 JS 2019-05 BugSwarm Tomassi et al. (2019) 1827/1264 Java/Python 2019-05 CPatMiner Nguyen et al. (2019) 17K Java 2019-05 ManySStuBs4J Karampatsis & Sutton (2020) 154K Java 2019-11 Refactory Hu et al. (2019) 1783 Python 2020-07 CoCoNut Lutellier et al. (2020) 24M Java, JS, C, Python 2020-10 Review4Repair Huq et al. (2022) 58021 Java 2020-11 BugsInPy Widyasari et al. (2020) 493 Python 2021-07 TFix Berabi et al. (2021) 105K JS 2021-08 Megadiff Monperrus et al. (2021) 663K Java 2022-01 SSB/TSSB Richter & Wehrheim (2022) 9M/3M Python 2022-10 FixJS Csuvik & Vidács (2022) 324K JS 2022-11 TypeBugs Oh & Oh (2022) 93 Python 2023-03 xCodeEval Khan et al. (2023) 4.7M 11 2023-04 RunBugRun Prenner & Robbes (2023) 450K C, C++, Java, Python, JS, Ruby, Go, PHP 2023-08 HumanEvalPack Muennighoff et al. (2023) 984 Python, JS, Go, Java, C++, Rust
Task Date Benchmark Source Size Language Code Summarization 2016-08 CODE-NN Iyer et al. (2016) 66K/32K C#/SQL 2017-07 unnamed Barone & Sennrich (2017) 150K Python 2018-05 DeepCom Hu et al. (2018a) 588K Java 2018-07 TL-CodeSum Hu et al. (2018b) 411K Java 2018-11 unnamed Wan et al. (2018) 109K Python 2019-02 unnamed LeClair et al. (2019) 2.1M Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M Go, JS, Python, PHP, Java, Ruby 2023-08 HumanEvalPack Muennighoff et al. (2023) 984 Python, JS, Go, Java, C++, Rust Code Completion 2013-05 GitHub Java Corpus Allamanis & Sutton (2013) 2.1M Java 2016-10 Py150 Raychev et al. (2016a) 150K Python 2016-10 JS150 Raychev et al. (2016a) 150K JS 2023-06 LCC Guo et al. (2023b) 360K Python, Java, C# Commit Message Generation 2017-03 unnamed Jiang & McMillan (2017) 509K Java 2017-04 CommitGen Loyola et al. (2017) 153K Python, C++, Java, JS 2017-08 CommitGen Jiang et al. (2017) 32K/75K Java 2018-09 NNGen Liu et al. (2018) 27K Java 2019-05 PtrGNCMsg Liu et al. (2019c) 64.9K Java 2019-08 CoDiSum Xu et al. (2019a) 90.7K Java 2019-12 ATOM Liu et al. (2022c) 160K Java 2021-05 CommitBERT Jung (2021) 346K Python, PHP, Go, Java, JS, Ruby 2021-07 MCMD Tao et al. (2021) 2.25M Java, C#, C++, Python, JS 2021-07 CoRec Wang et al. (2021b) 107K Java 2023-07 ExGroFi Wang et al. (2023a) 19263 Java 2023-08 CommitChronicle Eliseeva et al. (2023) 10.7M 20 Defect (Vulnerability) Detection 2018-01 CGD Li et al. (2018c) 62K C, C++ 2018-07 Draper VDISC Russell et al. (2018) 12.8M C, C++ 2018-07 SySeVR Li et al. (2018b) 15591 C, C++ 2018-04 unnamed Lin et al. (2018a) 32988 C, C++ 2019-02 unnamed Ponta et al. (2019) 624 Java 2019-09 Devign Zhou et al. (2019) 48687 C 2019-11 unnamed Lin et al. (2021) 170K C, C++ 2019-12 GREAT Hellendoorn et al. (2020) 2.8M Python 2020-01 MVD Zou et al. (2021) 182K C, C++ 2020-02 unnamed Lin et al. (2019) 1471 C 2020-09 ReVeal Chakraborty et al. (2022c) 18K C 2020-09 Big-Vul Fan et al. (2020) 265K C, C++ 2021-02 D2A Zheng et al. (2021) 1.3M C, C++ 2021-05 PyPIBugs Allamanis et al. (2021) 2374 Python 2021-07 CVEfixes Bhandari et al. (2021) 5495 27 2021-08 CrossVul Nikitopoulos et al. (2021) 27476 40+ 2023-04 DiverseVul Chen et al. (2023d) 349K C, C++ 2023-06 VulnPatchPairs Risse & Böhme (2023) 26K C 2023-11 VulBench Gao et al. (2023d) 455 C
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) 600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) 15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) 13948 2023-06 CMMLU Li et al. (2023a) 11528 2023-09 CodeApex Fu et al. (2023) 250 Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) 300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) 513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023b) 744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) 21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python
Task Date Benchmark Source Size Language Unit Test / Assertion Generation 2014-12 SF110 Fraser & Arcuri (2014) 24K Java 2020-09 Method2Test Tufano et al. (2021a) 781K Java 2021-08 ConTest Villmow et al. (2021) 365K Java Log Parsing 2018-11 LogHub (2018) Zhu et al. (2019); He et al. (2020) 379M 2023-08 LogHub (2023) Jiang et al. (2023c) 50.4M Repository-Level Coding 2023-03 RepoEval Zhang et al. (2023b) 1600/1600/373 Python 2023-06 RepoBench Liu et al. (2023i) 890K/9M/43K Python, Java 2023-06 PragmaticCode Agrawal et al. (2023) 880 Java 2023-06 Stack-Repo Shrivastava et al. (2023a) 816K Java 2023-09 CodePlan Bairi et al. (2023) 645/21 C#/Python 2023-10 SWE-Bench Jimenez et al. (2023) 2294 Python 2023-10 CrossCodeEval Ding et al. (2023) 9928 Python, Java, TypeScript, C#